OpenStudy (anonymous):

What is the solution to the equation 21^(x+3)=14? I need the answer in decimal form.

hartnn (hartnn):

this question belongs to logarithm topic, right ??

OpenStudy (anonymous):

right

OpenStudy (anonymous):

log(14)/log(21) = x+3 (log(14)/log(21)-3 = x

OpenStudy (anonymous):

(log(14)/log(21)) - 3 = x\ sorry forgot the parenthesis

hartnn (hartnn):

first step will be to take log on both sides :)

hartnn (hartnn):

then use the property \(\Large \log a^b = b\log a\)

OpenStudy (anonymous):

yeah, I got that, but all my answers are in approximate decimal form and I don't know how to get my equation to decimal form.

hartnn (hartnn):

so, whats the last step you're stuck at ? you got this ? \(x = \log 14/\log 21 -3 \) ?

OpenStudy (anonymous):

yeah, I just need it in decimal form

OpenStudy (nurali):

21^(x+3)=14 Put ln both sides ln21^(x+3)=ln14 Put down the exponent. (x+3)ln(21)=ln(14) divide both sides by ln(21) x+3=(ln14)/ln(21) subtract both side by 3 x=(ln14)/ln(21)-3 x=0.867-3 x=-2.13

hartnn (hartnn):

there you go ^ or you could have just used a calculator or even google :P

OpenStudy (anonymous):

wow, awesome! thank you guys so much!

OpenStudy (anonymous):

I like learning how to do things better than just getting the answer

hartnn (hartnn):

thats a pretty cool way of learning, i am glad :)

OpenStudy (nurali):

Anytime.